OpenStudy (anonymous):

Solve: y over negative 2 + 5 = 13

OpenStudy (unklerhaukus):

\[\frac{y}{-2}+5=13\]?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

take away five from both sides then multiply both sides by-1 you will have y=...

OpenStudy (anonymous):

idk mind telling me

OpenStudy (unklerhaukus):

\[\frac{y}{-2}+5-5=13-5\]\[\frac{y}{-2}=13-5\]\[-2\frac{y}{-2}=-2(13-5)\]\[\cdots\qquad=\qquad\cdots\]

OpenStudy (anonymous):

16/

OpenStudy (unklerhaukus):

nope

OpenStudy (anonymous):

-16

OpenStudy (unklerhaukus):

yep thats it you can even check in the original equation to be sure y over negative 2 + 5 = 13 -16 over negative 2 + 5 = 13 8+5=13